OpenStudy (anonymous):

Wanda filled a fish tank with water until the water level reached 4 in and then she added rocks to the tank and the water level rose to 7 inches. the tank is a rectangular prism with a base measuring 14in by 23 in. if Wanda added 85 rcks to the tank, what was the average volume of one of the rocks

Directrix (directrix):

The water in the fish tank was at height 4" before the rocks were added. Aftter the addition of rocks, the height of the water and rocks is 7" So, a 3" rise is height can be attibuted to the rocks alone. Find the volume of a rectangular prism within the tank, a prism with dimentions 14, 23, and 3. Volume of water rise from rocks = 14 * 23 * 3 = 966 cubic inches Divide that by the number of rocks (85) to get 966/85 = 11.364 cubic inches as the volume of the average rock added to the tank.